OpenStudy (anonymous):

D(h)=5e^(-0.32)h After 7 Hours there will be how many milligrams?

OpenStudy (anonymous):

I am guessing h = e? if so, D(7) = 5(7)^(-0.32) Solve for the exponent, then multiply.

OpenStudy (anonymous):

I am sorry! the h is after the (-0.32)

OpenStudy (anonymous):

Okay, D(7) = 5*e^(-.32)(7)

OpenStudy (anonymous):

ok how would I get that answer to be 0. something?

OpenStudy (anonymous):

What is e equal to?

OpenStudy (anonymous):

I am getting 25.42

OpenStudy (anonymous):

I think we can plug in e = 0 D(7) = 5*(0)^(-.32)(7) D(7) = 5*(0)*(7) D(7) = (0)*7 D(7) = 0